# When Text Embedding Meets Large Language Model: A Comprehensive Survey

### v3

@@ 5 Text Embedding Understanding with LLMs
LLMs have a powerful paraphrase capability that can be quickly aligned with and interpreted in a variety of already learned embedding spaces of image [227], item [228, 229] and concept [230]. Thus, for different purposes, existing work attempts to make an understanding of text embeddings with the help of LLMs. There are two embedding-related tasks in NLP: long context compression (ICC) and embedding inversion (EI).
@@
Embedding Source LLM Scale Form Paradigm Projector Task
PC [231] Soft Prompt GPT-2 300K X KL - PC
Gist [232] Soft Prompt LlaMa, FLAN-T5 130K (I, X, Y) KL - ICL
Gist-COCO [233] FLAN-T5 FLAN-T5 200K (I, X, Y) KL - ICL, RAG
HD-Gist [233] LlaMa LlaMa 50K (I, X, Y) KL - RAG
QGC [234] LlaMa2 LongChat, LlaMa2 300K (I, X, Y) LM+KL N-Layer MLP RAG
AutoCompressor [235] OPT, LlaMa OPT, LlaMa 2B X LM - PC
xRAG [135] SFR Mistral 955K (I, X, Y) LM→LM+KL 2-Layer MLP RAG
ReadOnce [236] BART BART 300K (I, X, Y) LM 1-Layer MLP RAG
ExtenToken [237] LlaMa2 LlaMa2 100K (I, X, Y) LM - RAG
UniICL [238] Vicuna Vicuna 10K (I, X, Y) LM 1-Layer MLP ICL
SCC [239] LlaMa LlaMa 500K (I, X, Y) LM - RAG
ICAE [136] LlaMa LlaMa 240K X AE+LM - PC
IC-Former [240] LlaMa2 LlaMa2 2.3M X AE→LM - PC
RCC [241] pythia, mamba pythia, mamba 5M X AE→LM - PC
PE-RANK [242] Jina-Embeddings, BGE-base Mistral 32M X AE→LM 2-Layer MLP IR
@@
5.1 Methods for Long Context Compression
Long Context Compression proposes that the input for conditioning a language model can be condensed into a smaller, specifically chosen set of words or dense vectors. These methods reduce context length, leading to faster inference speeds and smaller storage requirements. Based on the form of compressed context, this method can be classified into two categories: soft prompt 5.1.1 and context distillation 5.1.2. Long-context understanding has always been a significant issue in NLP. Some works [249, 250, 251, 252, 253, 254, 255] enhance the transformer’s ability to model long contexts by storing and utilizing past representations. [256] proposed compressing the raw activations of long contexts. These works primarily aim to expand the context of Transformers rather than compressing contexts, but their methods might also be adapted for context compression in the future.
5.1.1 Soft Prompt
Prompt tuning [257] introduces the trainable, parameterized soft prompt to adapt pre-trained language models to specific downstream tasks in natural language processing. In prompt compression, the general idea of this method is to align the soft prompt with the representation of the original long text, thereby enabling the replacement of the original long text with the soft prompt in practical applications. Different methods propose different optimization objectives to obtain sufficiently short soft prompts that can preserve the semantic meaning of the original text as much as possible. According to the training tasks, they can be divided into three categories: alignment, prediction, and restoration.
Alignment-Based Methods
The alignment method uses Kullback-Leibler (KL) divergence as the optimization objective. [231] train a small set of adaptable soft prompt weights to closely mimic the distribution induced by a larger, fixed hard prompt using KL divergence, resulting in a compact representation that retains essential information for guiding language model generation. Gist [232] adopts a special attention mask strategy in training to align gist tokens and task prompts. [258] creates hierarchical and dynamic gist tokens based on Gist for tool usage scenarios. [233] also uses KL divergence to align the gist representations obtained from the encoder with the original context. [234] uses queries to guide the context compression process and align the output by the LLM with the real answer using KL divergence.
Prediction-Based Methods
Restoration-Based Methods
5.1.2 Context Distillation
Selection-Based Methods
Generation-Based Methods
The generative method relies on appropriate text summarization techniques. [264] trains a small Language Model (LLM) by aligning embeddings of compressed context and original context while exploiting language model loss using compressed context in specific tasks. It ensures both the natural language format and the consistent utility of the original context. [265] exploits the text summarization technique to transform the original context into shorter and more coherent natural language summaries. [266] uses a pre-trained embedding model to extract sentence features and cluster similar sentences into text blocks. Then, text summarization technique is also used to obtain the final context. [267] scores each text block in the context and ultimately extracts context snippets as the compressed prompt. The advantage of this method is that it can produce natural language paragraphs understandable to humans, but it also comes with higher training costs. [268] leverages the prompt’s textual information to construct a graph, then utilizes a graph encoder to extract important elements from the graph to generate a summary.
@@ 6.1.1 False Negative detection
Under the embedding training framework of contrastive learning, the model aims to learn appropriate embeddings by pulling positive sample pairs closer while pushing negative sample pairs farther apart. The appearance of false negatives disrupts this learning process. Samples that originally belonged to the same category (positive sample relationship) are wrongly regarded as negative samples to participate in the training. This causes the model to receive incorrect "supervisory signals", resulting in the learned embeddings failing to accurately reflect the real semantic relationships among samples. An intuitive idea to solve the false negative problem is to correct the wrongly labeled samples through accurate annotations. However, whether it is the traditional manual annotation method or annotation with the help of large-scale language models (LLMs), both face the challenge of high costs. For manual annotation, in the face of massive data, it is almost an impossible task to expect to annotate thousands of samples for each query. Manual annotation is not only time-consuming and labor-intensive but also easily affected by subjective factors of annotators and issues such as the consistency of annotation standards. On the other hand, although LLMs have demonstrated powerful capabilities in many fields such as natural language processing, using them for annotation also comes with relatively high cost expenditures, including the cost of calling APIs, possible errors in generated results, and limitations on the accuracy of annotating domain-specific knowledge. Therefore, relying on large-scale and accurate annotations to completely eliminate false negatives has significant obstacles in reality.
@@
To address the false negative problem and the challenge of annotation costs, some studies have proposed methods such as improving the loss function or negative sample sampling strategies to attempt to alleviate this problem. For example, the peer loss method [277] designs a regularization term for the loss function, aiming to reduce the adverse impact of false negatives on the overall training gradient.  [278] believes that samples that are similar to positive samples but not similar to the query should be sampled, and such samples are more likely to be hard negatives rather than false negatives. However, these methods still have obvious limitations. On the one hand, they often rely heavily on the practical experience of researchers and lack a solid theoretical basis as a guarantee. This means that in different datasets and task scenarios, it is difficult to stably maintain the effectiveness of the methods, and there are doubts about their generalization ability. On the other hand, many improved loss functions have only been experimentally verified on a small range of data, and it cannot be proved whether they can truly and completely solve the false negative problem in a large-scale and diversified data environment. When applied to more complex and larger-scale practical tasks, these methods may not achieve the expected results, and the model training will still be interfered by false negatives, resulting in a significant reduction in the quality of embeddings and the performance of subsequent applications.
@@ 7.1.1 Cross-Lingual Text Embedding
The common backbones used by LLM-based embedders, such as Mistral and Qwen2, are pre-trained on large multi-lingual corpora, which leads us to expect higher-quality cross-language text embeddings using LLMs. UDEVER [196] shows the powerful embedding generalization capabilities of multi-lingual LLM, even when fine-tuned using only English text. Subsequent embedders based on Mistral or Qwen, such as GTE-Qwen2, GirtLM, and E5-Mistral, have achieved extraordinary performance on MTEB after contrastive learning on large-scale multi-lingual corpora. Specifically, E5-Mistral [157] synthesized a large-scale training corpus using LLM to enhance data diversity; GTE-Qwen2 [198] introduced weakly-supervised-supervised two-stage contrastive learning, which fully utilized the massive noisy web corpus. xVLM2Vec [311] extends cross-modal embeddings from MLLM-based embedders further in cross-language settings through self-knowledge distillation. Gemini-Embedding [312] integrates the experience of synthetic data and two-stage contrastive learning and, based on this, utilizes the technique of model merging [313] to further improve the performance of embedding on multi-lingual tasks.
@@ 7.1.2 Cross-Modal Text Embedding
In the cross-modal domain, joint language-vision embedding learning [334, 335, 336] has received the most attention. Its applications have gradually expanded from image-text retrieval [337, 338] to multiple tasks such as composed image retrieval [339, 340], multi-modal document retrieval [341] and multi-modal knowledge retrieval [342, 343], etc. Before the advent of multimodal large language models (MLLMs), ONE-PEACE [344] raised the number of parameters to 4B based on customizing the architecture, enabling a unified representation of images, text, and audio.
@@
Unlike text data, multimodal data is smaller-scale and more noisy, hindering the development of multi-modal Embedding. Recent studies have shown that using LLM to augment multi-modal data while employing good filtering will effectively improve the performance of multi-modal embedding across multiple tasks. For example, VISTA [345] generates high-quality instructions and captions with the powerful OpenAI LLM and Stable Diffusion, achieving substantial improvements on various cross-modal retrieval tasks. SPN [346] uses MLLMs and the pre-trained image encoder to synthesize both positive and negative examples for the composed image retrieval task, successfully improving the performance of multiple baseline models. To train the MLLM-based embedder, the larger scale instruction multi-modal dataset, such as MegaPairs [327] and VIRA [330], are synthesized with LLM-MLLM pipelines. MLLM-Embedder trained on these datasets show strong performance on uni-modal (even visual modal) and multi-modal tasks.
@@
As shown in Table 6, many attempts at MLLM-based embedders have been inspired by the successful experience of LLM-based embedders. For example, E5-V [321] follow the prompt style of PromptEOL [170] and fine-tune MLLMs with text contrastive learning only, firstly showing the potential of MLLMs as a multi-modal unified embedder. LLM2CLIP [322] trains a LLM-based embedder following LLM2Vec [168] and replaces the text encoder of CLIP with this LLM-based embedder, improving the performance of the CLIP model on multiple tasks effectively. MM-Embed [331] introduces the incremental text contrastive learning stage after traditional multi-modal contrastive learning, showing strong performance in both text and multi-modal embedding tasks. VladVA [326] uses next token prediction to increase the task difficulty for long text, preventing early gradient dissipation in the optimization process. In addition, many effective strategies in text embedding methods are followed, such as instruction tuning [323, 329], hard negative mining [328, 331], scaling negative number [333], etc. Except for MLLM-based embedders for universal multi-modal and uni-modal tasks, some MLLM-based embedders focus on the optimization of complex tasks in real scenarios. For example, DSE [324], UniSE-MLLM [330], and ColPali [332] focus on tasks related to document screenshot understanding. They leverage MLLM’s powerful comprehension capabilities to simplify the original complex embedding process while improving performance and efficiency. InstructCIR [325] focuses on composed image retrieval, which achieves better instruction-following capability by two-stage contrastive learning while progressive freezing the modules in the MLLM.
@@
As the number of MLLM-based embedders increased, evaluation benchmarks with different focuses were proposed, such as M-BEIR [336] for multi-modal retrieval, MVRB [330] for multi-modal document understanding, MMTB [329] for multi-modal universal embedding and UMRB [328] for both uni-modal and multi-modal universal embedding. However, we find that many works do not indicate whether their evaluation is a true zero-shot out-of-domain or not, and there was a risk of unfair performance comparison. This is the reason why we do not list “evaluation-setting” in Table 6 as we do in Table 4. We call on subsequent work to filter the training data strictly to prevent overlapping training-evaluation data, providing a fair performance benchmark for the community.
@@ 7.2 Task-Specific Text Embedding
For a long time, text embedding learning has been expected to obtain a universal embedding without making the downstream task visible. However, this setting is often difficult in practice, such as (1) In the unsupervised setting, contrastive learning brings the two views of data augmentation for the same instance closer together, i.e., learning an encoder that is insensitive to a particular data augmentation transform. However, this may result in valuable information for downstream tasks not being included in the embedding [347]; (2) In the supervised setting, undifferentiated joint training of different tasks will hurt performance [348, 67]; (3) In evaluation, the two texts can be evaluated from various perspectives and show different semantic similarities [349, 350];
@@
The universal encoders accept the text input $x$ and the downstream task $t$ and represent $x$ specific to task $t$, where $t$ can be an identifier or a paragraph description of the task requirements. According to where $t$ acts, we can categorize the methods into pre-processing (also called instruction-following embedding) and post-processing (also called equivariant embedding learning).
@@ 7.2.1 Instruction-Following Embedding
In instruction-following embedding learning, $F$ is divided into two parts: $F=f\circ p_{t}$, where $p_{t}$ is a prompt template related to task $t$. In the beginning, prompts are used to align the fine-tuning targets with the pre-trained targets of the PLMs [351]. Then, similar instruction tuning [352] is proposed to adjust LLMs to be more adaptable to the format of task-oriented conversations. Instruction-following embedding learning follows the idea of instruction tuning but models all tasks into a contrastive paradigm. TART [51] firstly validates the feasibility of this idea to information retrieval, while Instructor [67] fine-tuning GTR on 330 NLP tasks and found that including instructions would alleviate conflicts between tasks compared to traditional multi-task learning. In the zero-shot evaluation, the model fine-tuned with instructions shows good generalization to new instructions and datasets. The findings related to this technique are still iterating: LLM2Vec [168] add instruction only on evaluation and demonstrate its effectiveness; Inbedder [197] demonstrate that question-and-answer pairs can significantly improve a model’s ability to comply with instructions under a particular training method.

### v4

@@ 5 Text Embedding Understanding with LLMs
Large language models have a powerful paraphrase capability that can be quickly aligned with and interpreted in a variety of already learned embedding spaces of image [268], item [269, 270] and concept [271]. Thus, for different purposes, existing work attempts to make an understanding of text embeddings with the help of LLMs. There are two embedding-related tasks in NLP: long context compression (ICC) and embedding inversion (EI).
5.1 Long Context Compression
Long context compression proposes that the input for conditioning a language model can be condensed into a smaller, specifically chosen set of words or dense vectors. This reduces context length, leading to faster inference speeds and smaller storage requirements. Here, we distinguish long text compression from two similar research directions: (1) language models with memory mechanisms [272, 273, 274, 275, 276, 277, 278] and (2) key-value compression. The language model with memory mechanisms is designed to improve the in-context length of language modeling, and key-value compression is a technique proposed to improve the decoding efficiency of LLM based on key-value.
First, we can distinguish between long context compression and language modeling with memory mechanisms in design principle: Language modeling with memory mechanisms is for the native architectural design, and the memory module is part of the language model and requires full-parameter pre-training so that the model learns to read (comprehension) and write (update) the memory. In contrast, Long text compression is a post-adaptation of existing LLMs, which does not or slightly changes the parameters and architecture of the learned LLMs. The key to this task is to learn a compressor that can be aligned with the LLM input space.
Next, we can distinguish between long text compression and key-value compression in terms of the form of the input. Key-value compression aims to design an algorithm for only pruning the key-value cache, whose inputs are generated by the attention mechanism in LLM. On the contrary, the input of long text compression is text, although its output may be key-value embedding.
In addition, the methods of long text compression are usually categorized into two main groups based on the output form : (1) dense embeddings and (2) discrete tokens. We introduce the methods whose output are dense embeddings in Section 5.1.1 and focus on the methods whose output are discrete tokens in Section 5.1.2.
5.1.1 Soft Prompt
Prompt tuning [279] introduces the trainable, parameterized soft prompt to adapt pre-trained language models to specific downstream tasks in natural language processing. In prompt compression, the general idea of this method is to align the soft prompt with the representation of the original long text, thereby enabling the replacement of the original long text with the soft prompt in practical applications. Different methods propose different optimization objectives to obtain sufficiently short soft prompts that can preserve the semantic meaning of the original text as much as possible. According to the training tasks, these methods can be divided into three categories: alignment, prediction, and restoration.
Alignment-Based Methods
The alignment method uses Kullback-Leibler (KL) divergence as the optimization objective. PC [280] trains a small set of adaptable soft prompt weights to closely mimic the distribution induced by a larger, fixed hard prompt using KL divergence, resulting in a compact representation that retains essential information for guiding language model generation. Gist [281] adopts a special attention mask strategy in training to align gist tokens and task prompts. HD-Gist [282] creates hierarchical and dynamic gist tokens based on Gist for tool usage scenarios. Gist-COCO [283] also uses KL divergence to align the gist representations obtained from the encoder with the original context. QGC [284] uses queries to guide the context compression process and align the output by the LLM with the ground-truth answer using KL divergence.
Prediction-Based Methods
Restoration-Based Methods
5.1.2 Context Distillation
Extraction-Based Methods
Abstract-Based Methods
The generative method relies on appropriate text summarization techniques. [297] trains a small Language Model (LLM) by aligning compressed context and original context embeddings while exploiting language model loss using compressed context in specific tasks. It ensures both the natural language format and the consistent utility of the original context. [298] exploits the text summarization technique to transform the original context into shorter and more coherent natural language summaries. [299] uses a pre-trained embedding model to extract sentence features and cluster similar sentences into text blocks. Then, the text summarization technique is used to obtain the final context. [300] scores each text block in the context and extracts context snippets as the compressed prompt. The advantage of this method is that it can produce natural language paragraphs that are understandable to humans, but it also comes with higher training costs. [301] leverages the prompt’s textual information to construct a graph, then utilizes a graph encoder to extract important elements from the graph to generate a summary.
@@ 5.1.2 Context Distillation
Embedder Projector LLM Scale Form Paradigm Projector CR Task
@@ 6.1.1 False Negative detection
Under the embedding training framework of contrastive learning, the model aims to learn appropriate embeddings by pulling positive sample pairs closer while pushing negative sample pairs farther apart. The appearance of false negatives disrupts this learning process. Samples that initially belonged to the same category (positive sample relationship) are wrongly regarded as negative samples to participate in the training. This causes the model to receive incorrect "supervisory signals", resulting in the learned embeddings failing to accurately reflect the real semantic relationships among samples. An intuitive idea to solve the false negative problem is to correct the wrongly labeled samples through accurate annotations. However, whether it is the traditional manual annotation method or annotation with the help of large-scale language models (LLMs), both face the challenge of high costs. For manual annotation, in the face of massive data, it is almost an impossible task to expect to annotate thousands of samples for each query. Manual annotation is not only time-consuming and labor-intensive but also easily affected by subjective factors of annotators and issues such as the consistency of annotation standards. On the other hand, although LLMs have demonstrated powerful capabilities in many fields such as natural language processing, using them for annotation also comes with relatively high cost expenditures, including the cost of calling APIs, possible errors in generated results, and limitations on the accuracy of annotating domain-specific knowledge. Therefore, relying on large-scale and accurate annotations to eliminate false negatives has significant obstacles in reality.
@@
To address the false negative problem and the challenge of annotation costs, some studies have proposed methods such as improving the loss function or negative sample sampling strategies to attempt to alleviate this problem. For example, the peer loss method [316] designs a regularization term for the loss function, aiming to reduce the adverse impact of false negatives on the overall training gradient.  [317] believes that samples that are similar to positive samples but not similar to the query should be sampled, and such samples are more likely to be hard negatives rather than false negatives. However, these methods still have obvious limitations. On the one hand, they often rely heavily on the practical experience of researchers and lack a solid theoretical basis as a guarantee. This means that in different datasets and task scenarios, it is difficult to maintain the effectiveness of the methods stably, and there are doubts about their generalization ability. On the other hand, many improved loss functions have only been experimentally verified on a small range of data, and it cannot be proved whether they can truly and completely solve the false negative problem in a large-scale and diversified data environment. When applied to more complex and larger-scale practical tasks, these methods may not achieve the expected results, and the model training will still be interfered by false negatives, resulting in a significant reduction in the quality of embeddings and the performance of subsequent applications.
@@ 7.1.1 Cross-Lingual Text Embedding
The common backbones used by LLM-based embedders, such as Mistral and Qwen2, are pre-trained on large multi-lingual corpora, which leads us to expect higher-quality cross-language text embeddings using LLMs. UDEVER [195] shows the powerful embedding generalization capabilities of multi-lingual LLM, even when fine-tuned using only English text. Subsequent embedders based on Mistral or Qwen, such as GTE-Qwen2, GirtLM, and E5-Mistral, have achieved extraordinary performance on MTEB after contrastive learning on large-scale multi-lingual corpora. Specifically, E5-Mistral [157] synthesized a large-scale training corpus using LLM to enhance data diversity; GTE-Qwen2 [218] introduced weakly-supervised-supervised two-stage contrastive learning, which fully utilized the massive noisy web corpus. xVLM2Vec [350] extends cross-modal embeddings from MLLM-based embedders further in cross-language settings through self-knowledge distillation. Gemini-Embedding [216] integrates the experience of synthetic data and two-stage contrastive learning, and based on this, utilizes the model merging technique [255] to improve the embedding performance on multi-lingual tasks further.
@@ 7.1.2 Cross-Modal Text Embedding
In the cross-modal domain, joint language-vision embedding learning [371, 372, 373] has garnered the most attention. Its applications have gradually expanded from image-text retrieval [374, 375] to multiple tasks such as composed image retrieval [376, 377], multi-modal document retrieval [378] and multi-modal knowledge retrieval [379, 380], etc. Before the advent of multimodal large language models (MLLMs), ONE-PEACE [381] raised the number of parameters to 4B based on customizing the architecture, enabling a unified representation of images, text, and audio.
@@
Unlike text data, multimodal data is smaller-scale and noisier, hindering the development of multi-modal Embedding. Recent studies have shown that using LLM to augment multi-modal data while employing good filtering will effectively improve the performance of multi-modal embedding across multiple tasks. For example, VISTA [382] generates high-quality instructions and captions with the powerful OpenAI LLM and Stable Diffusion, achieving substantial improvements on various cross-modal retrieval tasks. SPN [383] uses MLLMs and the pre-trained image encoder to synthesize both positive and negative examples for the composed image retrieval task, successfully improving the performance of multiple baseline models. To train the MLLM-based embedder, the larger scale instruction multi-modal dataset, such as MegaPairs [364] and VIRA [367], are synthesized with LLM-MLLM pipelines. MLLM-Embedder trained on these datasets show strong performance on uni-modal (even visual modal) and multi-modal tasks.
@@
As shown in Table 5, many attempts at MLLM-based embedders have been inspired by the successful experience of LLM-based embedders. For example, E5-V [358] follow the prompt style of PromptEOL [169] and fine-tune MLLMs with text contrastive learning only, firstly showing the potential of MLLMs as a multi-modal unified embedder. LLM2CLIP [359] trains an LLM-based embedder following LLM2Vec [193] and replaces the text encoder of CLIP with this LLM-based embedder, improving the performance of the CLIP model on multiple tasks effectively. MM-Embed [368] introduces the incremental text contrastive learning stage after traditional multi-modal contrastive learning, showing strong performance in both text and multi-modal embedding tasks. VladVA [363] uses next token prediction to increase the task difficulty for long text, preventing early gradient dissipation in the optimization process. In addition, many effective strategies in text embedding methods are followed, such as instruction tuning [360, 366], hard negative mining [365, 368], scaling negative number [370], etc. Except for MLLM-based embedders for universal multi-modal and uni-modal tasks, some MLLM-based embedders focus on the optimization of complex tasks in real scenarios. For example, DSE [361], UniSE-MLLM [367], and ColPali [369] focus on tasks related to document screenshot understanding. They leverage MLLM’s powerful comprehension capabilities to simplify the original complex embedding process while improving performance and efficiency. InstructCIR [362] focuses on composed image retrieval, which achieves better instruction-following capability by two-stage contrastive learning while progressive freezing the modules in the MLLM.
@@
As the number of MLLM-based embedders increased, evaluation benchmarks with different focuses were proposed, such as M-BEIR [373] for multi-modal retrieval, MVRB [367] for multi-modal document understanding, MMTB [366] for multi-modal universal embedding and UMRB [365] for both uni-modal and multi-modal universal embedding. However, we find that many works do not indicate whether their evaluation is a true zero-shot out-of-domain or not, and there was a risk of unfair performance comparison. This is the reason why we do not list “evaluation-setting” in Table 5 as we do in Table 4. We encourage subsequent work to filter the training data strictly to prevent overlapping training and evaluation data, thereby providing a fair performance benchmark for the community.
@@ 7.2 Task-Specific Text Embedding
For a long time, text embedding learning has been expected to obtain a universal embedding without revealing the downstream task. However, this setting is often difficult in practice, such as (1) In the unsupervised setting, contrastive learning brings the two views of data augmentation for the same instance closer together, i.e., learning an encoder that is insensitive to a particular data augmentation transform. However, this may result in valuable information for downstream tasks not being included in the embedding [384]; (2) In the supervised setting, undifferentiated joint training of different tasks will hurt performance [385, 67]; (3) In evaluation, the two texts can be evaluated from various perspectives and show different semantic similarities [386, 387];
@@
The universal encoders accept the text input $x$ and the downstream task $t$, and represent $x$ specifically for task $t$, where $t$ can be an identifier or a paragraph description of the task requirements. According to where $t$ acts, we can categorize the methods into pre-processing (also called instruction-following embedding) and post-processing (also called equivariant embedding learning).
@@ 7.2.1 Instruction-Following Embedding
In instruction-following embedding learning, $F$ is divided into two parts: $F=f\circ p_{t}$, where $p_{t}$ is a prompt template related to task $t$. Initially, prompts are utilized to align the fine-tuning targets with the pre-trained targets of the PLMs [388]. Then, similar instruction tuning [389] is proposed to adjust LLMs to be more adaptable to the format of task-oriented conversations. Instruction-following embedding learning follows the idea of instruction tuning but models all tasks into a contrastive paradigm. TART [51] firstly validates the feasibility of this idea for information retrieval, while Instructor [67] fine-tuned GTR on 330 NLP tasks and found that including instructions would alleviate conflicts between tasks compared to traditional multi-task learning. In the zero-shot evaluation, the model fine-tuned with instructions shows good generalization to new instructions and datasets. The findings related to this technique are still iterating: LLM2Vec [193] adds instruction only on evaluation and demonstrates its effectiveness; Inbedder [196] demonstrates that question-and-answer pairs can significantly improve a model’s ability to comply with instructions under a particular training method.